OpenStudy (anonymous):

solve for p: |4p+3|/|p-5|=3

hero (hero):

\[\frac{|4p + 3|}{|p - 5|} = 3\] \[\left|\frac{4p + 3}{p - 5}\right| = 3\]

OpenStudy (anonymous):

i get that. and then you have a negative and positive 3 but how do you solve for p? multiply by the bottom half of the fraction?

OpenStudy (anonymous):

thank you

hero (hero):

Actually, for this all you have to do is just solve this \[\frac{4p + 3}{p - 5} = \pm 3\]

hero (hero):

\[\frac{4p + 3}{p - 5} = 3\] \[\frac{4p + 3}{p - 5} = -3\]